# REArtGS: Reconstructing and Generating Articulated Objects via 3D Gaussian Splatting with Geometric and Motion Constraints

### v3

@@ Abstract
Articulated objects, as prevalent entities in human life, their 3D representations play crucial roles across various applications. However, achieving both high-fidelity textured surface reconstruction and dynamic generation for articulated objects remains challenging for existing methods. In this paper, we present REArtGS, a novel framework that introduces additional geometric and motion constraints to 3D Gaussian primitives, enabling high-quality textured surface reconstruction and generation for articulated objects. Specifically, given multi-view RGB images of arbitrary two states of articulated objects, we first introduce an unbiased Signed Distance Field (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we establish deformable fields for 3D Gaussians constrained by the kinematic structures of articulated objects, achieving unsupervised generation of surface meshes in unseen states. Extensive experiments on both synthetic and real datasets demonstrate our approach achieves high-quality textured surface reconstruction for given states, and enables high-fidelity surface generation for unseen states. Codes will be released after acceptance and the project website is at https://sites.google.com/view/reartgs/home.
@@ 1 Introduction
Articulated objects are ubiquitous in our daily lives. Modeling articulated objects, i.e. mesh reconstruction and generation, holds significant importance in many fields of computer vision and robotics fields including virtual and augmented reality [17, 33], object manipulation [37, 29] and human-object interaction [34, 13]. Currently, the surface and shape reconstruction for articulated objects is a non-trivial task due to the challenges: firstly, articulated objects exhibit complicated and diverse geometric structures with a wide range of scales. Secondly, articulated objects possess varied kinematic structures, and the static surface meshes generated from vanilla 3D reconstruction methods fail to meet practical interaction requirements.
Under this circumstance, PARIS [15] attempts using multi-view images from two states to dynamically reconstruct articulated objects with neural implicit radiance fields. Nevertheless, we make observation that PARIS holds two problems leading to difficulty in reconstructing high-quality surface meshes: (1) It relies on neural implicit representation to reconstruct objects, lacking constraints on geometric learning, which leads to shape-radiance ambiguity and additional errors when extracting surface meshes from the learned entangled neural representations. (2) PARIS fails to generate texture-rich meshes for articulated objects, while texture is significant for bridging sim-to-real gap.
@@
In recent years, 3D Gaussian Splatting (3DGS) [11] achieves realistic and real-time novel view synthesis through explicit representation with 3D Gaussian primitives. Some subsequent works have expanded 3DGS to surface reconstruction [39, 7, 1] and dynamic reconstruction  [35, 14, 4]. However, existing 3DGS-based surface reconstruction methods typically suffer from insufficient geometric constraints [39] or impose constraints by restricting the shapes of Gaussian primitives  [7, 6], leading to artifacts or inaccurate recovery of detail structures. On the other hand, current dynamic reconstruction methods via 3DGS tend to input motion time and spatial positions into neural networks to obtain the deformed positions of Gaussian primitives [35, 4]. Consequently, these methods require continuous supervision throughout the entire motion, which limits their ability to generate surface meshes in unseen states. In general, introducing 3DGS as a ready-to-use technique into articulated object surface reconstruction and generation is still not feasible and remains a challenging task.
In this paper, to address the aforementioned issues and make full advantage of 3DGS, we propose REArtGS, a novel approach that Reconstructs and gEnerates high-quality texture-rich mesh surfaces for Articulated objects via 3D Gaussian Splatting, only taking multi-view RGB images from two arbitrary states. We firstly introduce an unbiased Signed Distance Function (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we employ the optimized Gaussian primitives as an accurate geometry initialization for dynamic surface generation. In addition, we leverage the kinematic structures of articulated objects to model time-continuous Gaussian deformable fields in an unsupervised fashion. Our REArtGS is evaluated on PartNet-Mobility [31] and AKB-48 [16] object repositories, ranging from synthetic to real-world data. Extensive experiments demonstrate that our REArtGS outperforms existing state-of-the-art methods in articulated object surface reconstruction and generation tasks.
@@ 2.1 Articulated Object Shape Reconstruction
Object surface reconstruction is a well-established problem for understanding the full geometric shape of objects. Some works introduce to encode continuous functions that model the objects using Signed Distance [23], radiance [20, 26, 28] and occupancy [19]. Going beyond static scenes with rigid objects, A-SDF [21] extends the DeepSDF to learn a latent code for modeling an articulated object by both shape and motion. Following up on A-SDF, CenterSnap [8] designs a point cloud decoder for shape reconstruction, and Shapo [9] fully utilizes shape prior for latent space learning. To generalize these works into the real world, AKBNet [16] integrates the shape reconstruction module into the object knowledge modeling along with part segmentation and pose estimation. To achieve both geometry and motion analysis in a single forward, Ditto [10] and REACTO [27] propose to generate shapes at unseen states from pair observations. Currently, PARIS [15] alleviate the 3D data requirement limitation and success to reconstruct shape surface with only RGB images. But it still suffer from several challenges such as lacking constraints modeling and failing in texture generation. Therefore, our REArtGS aims to address these issues, and proposes geometric and motion constrained 3DGS for this task.
@@ 2.2 Surface Reconstruction with 3DGS
3D Gaussain Splatting has become increasingly popular technique for surface reconstruction in recent years [11, 12, 38, 25]. GOF [39] establishes opacity fields of 3D Gaussians using ray-tracing-based rendering, and extracts the surface meshes by the opacity level set. 2DGS [7] and PGSR [1] seek to transform 3D Gaussians into 2D flat representation, obtaining accurate normal distribution. Although improving the surface reconstruction quality, these methods still lack more reasonable geometry constraints. Several researchers attempt to integrate SDF representation with 3D Gaussians [36, 2, 32, 3], but these approaches commonly use SDF to regularize normals and guide the pruning of 3D Gaussians, without substantial optimization of Gaussian opacity fields. Meanwhile, some works conduct dynamic reconstruction using 3DGS. Deformable 3DGS proposes a deformation field to reconstruct dynamic scenes with 3D Gaussian primitives. 4DGS [4] leverages a spatial-temporal structure encoder and a multi-head Gaussian deformation decoder to derive the deformed 3D Gaussians at a given time. However, these methods rely on complete supervision of the motion process, limiting their capacity for generating unseen states. The similar issues also occur in DGMesh [14] and REACTO [27]. Additionally, these methods lack an adequate initialization for the canonical geometry, leading to a decline in surface quality when directly conducting dynamic rendering.
@@ 3.1 Reconstruction with Unbiased SDF Guidance
Following Neus [30], we can derive the opacity $\bm{\hat{\sigma}}$ for the center $\mathbf{x}_{i}$ of Gaussian $G_{i}$ from learned SDF values:
@@
[…]
@@
$\mathcal{L}_{unbias}=\left\|f(\mathbf{o}+t^{*}\mathbf{r})\right\|_{2}^{2}$ (8)
where $\mathbf{o}$, $\mathbf{r}$ are the camera center and incident ray direction. Through the unbiased regularization, we encourage that when the rendering contribution of a Gaussian primitive reaches its maximum value, the corresponding spatial position is close to the scene surface, as shown in Fig. 3. In this manner, the SDF representation is able to regularize the opacity fields without bias and facilitate more reasonable distribution of Gaussian primitives over scene surface. We provide more detailed elaboration and proof in the supplementary materials.
@@ 3.2 Mesh Generation with Motion Constraints
Due to the explicit representation of Gaussian primitives, we can acquire accurate dense point clouds from reconstruction, serving as initial geometry for mesh generation. Then we establish time-continuous deformable fields for 3D Gaussians to infer the deformed position $\mathbf{x}_{s}$ at state $s$. However, it is challenging to optimize the deformable fields for Gaussian primitives only with the supervision of two states. Accordingly, we exploit the kinematic structures of articulated objects to constrain the Gaussian deformation field. We first assume that the articulated motion only includes rotation and translation following PARIS. For a rotation joint, our learning objective is to determine its rotation axis and rotation angle. Specifically, we take the pivot point $\mathbf{o}_{r}\in\mathbb{R}^{3}$ of the rotation axis and a normalized quaternion $\mathbf{q}\in\mathbb{R}^{4}$ as learnable parameters, where the quaternion can be decoupled into the rotation axis $\mathbf{a}\in\mathbb{R}^{3}$ and rotation angle $\theta$. For a prismatic joint, our learning objective is to determine their translation directions and translation distances. Similarly, we take the unit vector $\mathbf{d}\in\mathbb{R}^{3}$ representing the translation direction and the translation distance $m$ as learnable parameters.
@@
where $s\in\left[0,1\right]$ and $\theta\in\left[-\pi/2,\pi/2\right]$. The angle-bounded parameterization is able to prevent the singularity in exponential coordinates when $\left\|\theta\right\|>\pi$. Using Rodrigues’ rotation formula [24], the deformed position $\mathbf{x}_{s}$ can be derived as following:
@@
For the prismatic motion $m\cdot\mathbf{d}\in\mathbb{R}^{3}$, we take advantage of the vector space structure of Euclidean geometry by setting $s^{*}=0$. The deformed position $\mathbf{x}_{s}$ can be naturally acquired by linear interpolation without additional constraints, formulated as:
@@ 3.3 Optimization and Textured Mesh Extraction
$\mathcal{L}_{normal}=\frac{1}{N}\sum_{i}^{N}(1-\frac{\left|\mathbf{n}_{i}\cdot\nabla f(\mathbf{x}_{i})\right|}{\left\|\mathbf{n}_{i}\right\|\cdot\left\|f(\mathbf{x}_{i})\right\|})$ (13)
@@
where $R$ is the rotation matrix encoded by Gaussian’s quaternion and $S$ is the scaling matrix. We also employ the Eikonal regularization term [5] to encourage SDF gradient direction perpendicular to the surface, represented as:
$\mathcal{L}_{eik}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla f(\mathbf{x}_{i}))\right\|_{2}-1\right)^{2}$ (15)
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{unbias}+\lambda_{2}\mathcal{L}_{normal}+\lambda_{3}\mathcal{L}_{eik}+\lambda_{3}\mathcal{L}_{d}$ (16)
where $\lambda$ is the weight of regularization and $\mathcal{L}_{d}$ is the depth distortion loss following [7]. $\mathcal{L}_{c}$ is defined as:
$\mathcal{L}_{c}=\lambda\mathcal{L}_{D-SSIM}+(1-\lambda)\mathcal{L}_{1}$ (17)
where $\mathcal{L}_{D-SSIM}$ is proposed in [11] and $\mathcal{L}_{1}$ is the L1 norm of the pixel loss.
Once the optimization converges, we adopt the TSDF fusion algorithm[22] to extract the textured mesh from Gaussian primitives. We utilize the $\alpha$-blending rendering pipeline to obtain depth, opacity and RGB renderings for training views. Then we integrate these images into a voxel block grid (VBG) and extract a triangle mesh from the VBG. The textures of surface meshes can be efficiently obtained leveraging the spherical harmonics and opacity of Gaussian primitives. Please refer to our supplementary materials for a more detailed illustration.
Figure 4: The qualitative result of surface reconstruction on PartNet-Mobility dataset. We show both textured and non-textured meshes for best comparison.
@@
CD(ws) $\downarrow$ A-SDF [21] 14.19 7.14 10.61 13.71 40.85 12.50 3.31 2.11 21.37 22.57 14.84
Ditto [10] 2.38 2.09 1.70 2.16 6.80 7.29 42.04 0.31 2.51 3.91 7.19
PARIS [15] 0.96 1.80 0.30 2.68 0.42 18.31 0.46 0.25 6.07 8.12 3.94
GOF [39] 1.51 8.09 0.36 1.51 0.53 17.35 0.89 0.84 19.81 10.04 6.09
ArtGS [18] 2.77 1.36 0.75 2.01 0.41 20.59 0.63 0.99 9.01 9.00 4.75
@@
CD(rs) $\downarrow$ A-SDF [23] 2.140 2.478 1.805 1.801 2.080 1.590 5.630 2.418 2.751 4.077 2.677
Ditto [10] 2.874 3.049 2.926 1.550 0.925 1.428 5.792 1.184 1.296 2.837 2.386
PARIS [15] 2.077 3.812 2.807 0.370 1.209 3.869 2.855 1.025 2.873 2.903 2.380
GOF [39] 2.510 1.736 2.304 2.230 2.787 5.725 2.111 0.961 3.080 2.812 2.626
ArtGS [18] 2.949 1.704 2.438 0.725 0.615 2.874 2.031 1.182 1.148 1.157 1.682
@@
F1 $\uparrow$ A-SDF [23] 0.041 0.035 0.094 0.019 0.053 0.046 0.224 0.001 0.010 0.015 0.054
Ditto [10] 0.197 0.181 0.275 0.114 0.352 0.059 0.107 0.366 0.052 0.030 0.173
PARIS [15] 0.240 0.151 0.343 0.091 0.429 0.024 0.396 0.533 0.031 0.033 0.227
GOF [39] 0.217 0.189 0.614 0.173 0.480 0.087 0.386 0.303 0.049 0.036 0.253
ArtGS [18] 0.251 0.209 0.438 0.118 0.447 0.053 0.408 0.294 0.056 0.042 0.232
@@
EMD $\downarrow$ A-SDF [23] 0.755 1.113 0.952 0.945 1.020 0.982 1.763 1.039 1.174 1.258 1.100
Ditto [10] 1.724 1.308 1.212 0.619 0.935 0.841 1.996 0.417 0.852 0.970 1.087
PARIS [15] 1.197 1.381 0.637 0.426 0.778 1.660 1.823 0.706 1.200 1.046 1.085
GOF [39] 1.643 1.815 1.075 1.055 1.181 1.692 1.237 0.394 1.241 0.959 1.229
ArtGS [18] 1.921 0.677 0.934 0.637 0.506 1.154 1.131 0.502 1.006 0.787 0.926
@@
CD(ws) $\downarrow$ A-SDF [21] 112.30 34.01 94.11 33.46 56.73 55.82 3.78 27.64 29.60 25.30 47.28
PARIS [15] 2.09 14.51 12.81 2.64 10.42 18.30 0.80 3.25 10.89 5.89 8.16
D-3DGS [35] 98.90 16.79 91.84 48.78 30.10 30.93 1.262 65.11 22.39 18.79 42.49
ArtGS [18] 2.34 1.72 0.68 2.05 0.52 20.59 0.63 1.00 8.78 10.97 4.92
@@
CD(rs) $\downarrow$ A-SDF [21] 8.992 1.806 1.822 5.166 5.668 7.050 5.991 2.710 5.245 7.794 5.224
PARIS [15] 2.113 0.417 4.704 0.408 1.216 0.857 4.442 1.354 2.793 4.332 2.264
D-3DGS [35] 9.589 1.544 0.891 0.465 1.147 3.218 0.168 0.449 3.773 3.100 2.434
ArtGS [18] 4.924 0.670 0.381 0.471 1.622 2.670 1.455 1.269 3.912 0.562 1.794
@@
F1 $\uparrow$ A-SDF [21] 0.007 0.034 0.056 0.012 0.021 0.023 0.168 0.001 0.010 0.004 0.034
PARIS [15] 0.221 0.151 0.319 0.091 0.423 0.024 0.421 0.530 0.031 0.032 0.224
D-3DGS [35] 0.064 0.180 0.125 0.073 0.168 0.195 0.233 0.131 0.022 0.023 0.121
ArtGS [18] 0.264 0.255 0.454 0.125 0.460 0.055 0.426 0.292 0.062 0.045 0.244
@@
EMD $\downarrow$ A-SDF [21] 2.123 0.951 0.955 1.606 1.684 1.878 1.731 1.115 1.620 1.837 1.559
PARIS [15] 2.454 0.457 1.533 0.451 0.780 0.655 1.491 0.770 1.183 1.354 1.113
D-3DGS [35] 3.224 0.612 0.665 0.481 0.758 1.269 0.291 0.436 1.374 1.206 1.032
ArtGS [18] 2.207 0.562 0.412 0.563 0.525 1.171 0.631 0.724 1.297 0.750 0.884
@@
Figure 6: The qualitative results of surface reconstruction on real-world AKB-48 [16] repository. The multi-view images are obtained from ReArtVal [17] dataset. We show both textured and non-textured meshes for best view.

### v4

@@ Abstract
Articulated objects, as prevalent entities in human life, their 3D representations play crucial roles across various applications. However, achieving both high-fidelity textured surface reconstruction and dynamic generation for articulated objects remains challenging for existing methods. In this paper, we present REArtGS, a novel framework that introduces additional geometric and motion constraints to 3D Gaussian primitives, enabling realistic surface reconstruction and generation for articulated objects. Specifically, given multi-view RGB images of arbitrary two states of articulated objects, we first introduce an unbiased Signed Distance Field (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we establish deformable fields for 3D Gaussians constrained by the kinematic structures of articulated objects, achieving unsupervised generation of surface meshes in unseen states. Extensive experiments on both synthetic and real datasets demonstrate our approach achieves high-quality textured surface reconstruction for given states, and enables high-fidelity surface generation for unseen states. Project site: https://sites.google.com/view/reartgs/home.
@@ 1 Introduction
Articulated objects are ubiquitous in our daily lives. Modeling articulated objects, i.e. mesh reconstruction and generation, holds significant importance in many fields of computer vision and robotics fields including virtual and augmented reality liu2022toward ; xue2021omad , object manipulation yu2024gamma ; wang2024rpmart and human-object interaction yang2022oakink ; li2024semgrasp . Currently, the surface and shape reconstruction for articulated objects is a non-trivial task due to the challenges: firstly, articulated objects exhibit complicated and diverse geometric structures with a wide range of scales. Secondly, articulated objects possess varied kinematic structures, and the static surface meshes generated from vanilla 3D reconstruction methods fail to meet practical interaction requirements. Under this circumstance, PARIS liu2023paris attempts to use multi-view images from two states for articulated object dynamically reconstruction with neural implicit radiance fields. Nevertheless, PARIS lacks geometric constraints, leading to shape-radiance ambiguity and additional errors for motion analysis.
@@
In recent years, 3D Gaussian Splatting (3DGS) kerbl20233d achieves realistic and real-time novel view synthesis through explicit representation with 3D Gaussian primitives. Some subsequent works have expanded 3DGS to surface reconstruction yu2024gaussian ; huang20242d ; chen2024pgsr and dynamic reconstruction  yang2024deformable ; liu2024dynamic ; duan20244d . However, existing 3DGS-based surface reconstruction methods typically suffer from insufficient geometric constraints yu2024gaussian or impose constraints by restricting the shapes of Gaussian primitives  huang20242d ; guedon2024sugar , leading to noisy surface reconstruction results. On the other hand, current dynamic reconstruction methods via 3DGS tend to input motion time and spatial positions into neural networks to obtain the deformed positions of Gaussian primitives yang2024deformable ; duan20244d . Consequently, these methods require continuous supervision throughout the entire motion, which limits their ability to generate surface meshes in unseen states. In general, introducing 3DGS as a ready-to-use technique into articulated object surface reconstruction and generation is still not feasible and remains a challenging task.
In this paper, to address the aforementioned issues and make full advantage of 3DGS, we propose REArtGS, a novel approach that Reconstructs and gEnerates high-quality texture-rich mesh surfaces for Articulated objects via 3D Gaussian Splatting, only taking multi-view RGB images from two arbitrary states. Specifically, unlike most 3DGS methods based on the opacity field, which may lead to noisy surface extraction due to the non-strict linearity of the opacity field, we first introduce the Signed Distance Function (SDF) field to facilitate geometry learning. Then we propose a geometry constraint which encourages the SDF values of the Gaussian primitives approach zero when their opacity values reach the maximum , achieving unbiased guidance of the SDF field over the opacity field. In this way, we can take advantage of the strong linearity of the SDF field to explicitly establish a connection between Gaussian opacity field and the scene surface, which significantly enhances geometric learning and reduces artifacts.
Subsequently, we employ the optimized Gaussian primitives as an accurate geometry initialization for dynamic surface generation. Due to the absence of intermediate states, we leverage the kinematic structures of articulated objects to model time-continuous Gaussian deformable fields in an unsupervised fashion, and constrain the deformable fields using learnable motion parameters. Concretely, we propose a heuristic method for unsupervised part segmentation and formulate the deformable fields of dynamic parts via the motion parameters. Our REArtGS is evaluated on PartNet-Mobility xiang2020sapien and AKB-48 liu2022akb object repositories, ranging from synthetic to real-world data. Extensive experiments demonstrate that our REArtGS outperforms existing state-of-the-art methods in articulated object surface reconstruction and generation tasks, as shown in Fig. 1.
@@ 2.1 Articulated Object Shape Reconstruction
Object surface reconstruction is a well-established problem for understanding the full geometric shape of objects. Some works introduce to encode continuous functions that model the objects using Signed Distance park2019deepsdf ; mu2021sdf , radiance mildenhall2021nerf ; rakotosaona2024nerfmeshing ; tang2023delicate and occupancy mescheder2019occupancy . To achieve both geometry and motion analysis in a single forward, Ditto jiang2022ditto and REACTO song2024reacto propose to generate shapes at unseen states from pair observations. Specifically, PARIS liu2023paris alleviates the 3D data requirement limitation and succeeds in reconstructing shape surface with only RGB images. Most recently, ArticulatedGS guo2025articulatedgs and ArtGS liu2025building introduce 3DGS to achieve both the reconstruction and motion estimation of articulated objects. However, these methods still suffer from insufficient geometry constraints, which may lead to noisy surface mesh outputs. Therefore, our REArtGS aims to address these issues, and proposes geometric and motion constrained 3DGS for this task.
@@ 2.2 Surface Reconstruction with 3DGS
3D Gaussain Splatting has become increasingly popular technique for surface reconstruction in recent years kerbl20233d ; li2024loopgaussian ; yu2024mip ; qin2024langsplat . GOF yu2024gaussian establishes opacity fields of 3D Gaussians using ray-tracing-based rendering, and extracts the surface meshes by the opacity level set. 2DGS huang20242d and PGSR chen2024pgsr seek to transform 3D Gaussians into 2D flat representation, obtaining accurate normal distribution. Although improving the surface reconstruction quality, these methods still lack more reasonable geometry constraints. Several researchers attempt to integrate SDF representation with 3D Gaussians yu2024gsdf ; chen2023neusg ; xiang2024gaussianroom ; dai2024high , but these approaches commonly use SDF to regularize normals and guide the pruning of 3D Gaussians, without substantial optimization of Gaussian opacity fields. Meanwhile, some works conduct dynamic reconstruction using 3DGS. Deformable 3DGS proposes a deformation field to reconstruct dynamic scenes with 3D Gaussian primitives. 4DGS duan20244d leverages a spatial-temporal structure encoder and a multi-head Gaussian deformation decoder to derive the deformed 3D Gaussians at a given time. However, these methods rely on complete supervision of the motion process, limiting their capacity for generating unseen states. The similar issues also occur in DGMesh liu2024dynamic and REACTO song2024reacto . We provide more analysis and comparison of the related works in Appendix.
@@ 3.1 Reconstruction with Unbiased SDF Guidance
Following Neus wang2021neus , we can derive the opacity $\bm{\hat{\sigma}}$ for the center $\mathbf{x}_{i}$ of Gaussian $G_{i}$ from learned SDF values:
@@
$t_{\text{bias}}=\left\|\underset{t}{\text{argmax}}(\bm{\hat{\sigma}})-\underset{t}{\text{argmax}}(\bm{\varepsilon)}\right\|$ (7)
@@
$\mathcal{L}_{\text{unbias}}=\left\|f(\mathbf{o}+t^{*}\mathbf{r})\right\|_{2}^{2}$ (8)
where $\mathbf{o}$, $\mathbf{r}$ are the camera center and incident ray direction. Through the unbiased regularization, we encourage that when the rendering contribution of a Gaussian primitive reaches its maximum value, the corresponding spatial position is close to the scene surface, as shown in Fig. 3. In this manner, the SDF representation is able to regularize the opacity fields without bias and facilitate more reasonable distribution of Gaussian primitives over scene surface. We provide more detailed elaboration and proof in the Appendix.
@@ 3.2 Mesh Generation with Motion Constraints
Due to the explicit representation of Gaussian primitives, we can acquire accurate dense point clouds from reconstruction, serving as initial geometry for mesh generation. Then we establish time-continuous deformable fields for 3D Gaussians to infer the deformed position $\mathbf{x}_{s}$ at state $s$. However, it is challenging to optimize the deformable fields for Gaussian primitives only with the supervision of two states. Accordingly, we exploit the kinematic structures of articulated objects to constrain the Gaussian deformation field. We first assume that the articulated motion only includes rotation and translation following PARIS. Then we employ a heuristic method to efficiently predict the joint type through movement trend of points, which is elaborated in the Appendix. For rotation joints, our learning objective is to determine their rotation axis and rotation angle. Specifically, we take the pivot point $\mathbf{o}_{r}\in\mathbb{R}^{3}$ of the rotation axis and a normalized quaternion $\mathbf{q}\in\mathbb{R}^{4}$ as learnable parameters, where the quaternion can be decoupled into the rotation axis $\mathbf{a}\in\mathbb{R}^{3}$ and rotation angle $\theta$. For prismatic joints, our learning objective is to determine their translation directions $\mathbf{d}$ and translation distances $m$. Similarly, we take the unit vector $\mathbf{d}\in\mathbb{R}^{3}$ representing the translation direction and the translation distance $m$ as learnable parameters.
@@
where $s\in\left[0,1\right]$ and $\theta\in\left[-\pi/2,\pi/2\right]$. The angle-bounded parameterization is able to prevent the singularity in exponential coordinates when $\left\|\theta\right\|>\pi$. Using Rodrigues’ rotation formula pina2011rotations , the deformed position $\mathbf{x}_{s}$ can be derived as following:
@@
For the prismatic motion $m\cdot\mathbf{d}\in\mathbb{R}^{3}$, we take advantage of the vector space structure of Euclidean geometry by setting $s^{*}=0$. The deformed position $\mathbf{x}_{s}$ can be naturally acquired by linear interpolation, formulated as:
@@
Moreover, our method can be conveniently transferred to multi-part articulated objects, through sequentially learning the segmentation masks and motions of each individual part. Please refer to our Appendix for more details.
@@ 3.3 Optimization and Textured Mesh Extraction
Figure 4: The qualitative result of surface reconstruction on PartNet-Mobility dataset. We show both textured and non-textured meshes for the best comparison.
@@
$\mathcal{L}_{\text{normal}}=\frac{1}{N}\sum_{i}^{N}(1-\frac{\left|\mathbf{n}_{i}\cdot\nabla f(\mathbf{x}_{i})\right|}{\left\|\mathbf{n}_{i}\right\|\cdot\left\|f(\mathbf{x}_{i})\right\|})$ (13)
@@
where $R$ is the rotation matrix encoded by Gaussian’s quaternion and $S$ is the scaling matrix. We also employ the Eikonal regularization term gropp2020implicit to encourage SDF gradient direction perpendicular to the surface, represented as:
$\mathcal{L}_{\text{eik}}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla f(\mathbf{x}_{i}))\right\|_{2}-1\right)^{2}$ (15)
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{\text{unbias}}+\lambda_{2}\mathcal{L}_{\text{normal}}+\lambda_{3}\mathcal{L}_{\text{eik}}+\lambda_{3}\mathcal{L}_{d}$ (16)
where $\lambda$ is the weight of regularization and $\mathcal{L}_{d}$ is the depth distortion loss following huang20242d . $\mathcal{L}_{c}$ is defined as:
$\mathcal{L}_{c}=\lambda\mathcal{L}_{\text{D-SSIM}}+(1-\lambda)\mathcal{L}_{1}$ (17)
where $\mathcal{L}_{\text{D-SSIM}}$ is proposed in kerbl20233d and $\mathcal{L}_{1}$ is the L1 norm of the pixel loss.
Once the optimization converges, we adopt the TSDF fusion algorithmnewcombe2011kinectfusion to extract the textured mesh from Gaussian primitives. We utilize the $\alpha$-blending rendering pipeline to obtain depth, opacity and RGB renderings for training views. Then we integrate these images into a voxel block grid (VBG) and extract a triangle mesh from the VBG. The textures of surface meshes can be efficiently obtained leveraging the spherical harmonics and opacity of Gaussian primitives. Please refer to our Appendix for a more detailed illustration.
@@
CD(ws) $\downarrow$ A-SDF mu2021sdf 14.19 7.14 10.61 13.71 40.85 12.50 3.31 2.11 21.37 22.57 14.84
Ditto jiang2022ditto 2.38 2.09 1.70 2.16 6.80 7.29 42.04 0.31 2.51 3.91 7.19
PARIS liu2023paris 0.96 1.80 0.30 2.68 0.42 18.31 0.46 0.25 6.07 8.12 3.94
GOF yu2024gaussian 1.51 8.09 0.36 1.51 0.53 17.35 0.89 0.84 19.81 10.04 6.09
ArtGS liu2025building 2.77 1.36 0.75 2.01 0.41 20.59 0.63 0.99 9.01 9.00 4.75
@@
CD(rs) $\downarrow$ A-SDF park2019deepsdf 2.140 2.478 1.805 1.801 2.080 1.590 5.630 2.418 2.751 4.077 2.677
Ditto jiang2022ditto 2.874 3.049 2.926 1.550 0.925 1.428 5.792 1.184 1.296 2.837 2.386
PARIS liu2023paris 2.077 3.812 2.807 0.370 1.209 3.869 2.855 1.025 2.873 2.903 2.380
GOF yu2024gaussian 2.510 1.736 2.304 2.230 2.787 5.725 2.111 0.961 3.080 2.812 2.626
ArtGS liu2025building 2.949 1.704 2.438 0.725 0.615 2.874 2.031 1.182 1.148 1.157 1.682
@@
F1 $\uparrow$ A-SDF park2019deepsdf 0.041 0.035 0.094 0.019 0.053 0.046 0.224 0.001 0.010 0.015 0.054
Ditto jiang2022ditto 0.197 0.181 0.275 0.114 0.352 0.059 0.107 0.366 0.052 0.030 0.173
PARIS liu2023paris 0.240 0.151 0.343 0.091 0.429 0.024 0.396 0.533 0.031 0.033 0.227
GOF yu2024gaussian 0.217 0.189 0.614 0.173 0.480 0.087 0.386 0.303 0.049 0.036 0.253
ArtGS liu2025building 0.251 0.209 0.438 0.118 0.447 0.053 0.408 0.294 0.056 0.042 0.232
@@
EMD $\downarrow$ A-SDF park2019deepsdf 0.755 1.113 0.952 0.945 1.020 0.982 1.763 1.039 1.174 1.258 1.100
Ditto jiang2022ditto 1.724 1.308 1.212 0.619 0.935 0.841 1.996 0.417 0.852 0.970 1.087
PARIS liu2023paris 1.197 1.381 0.637 0.426 0.778 1.660 1.823 0.706 1.200 1.046 1.085
GOF yu2024gaussian 1.643 1.815 1.075 1.055 1.181 1.692 1.237 0.394 1.241 0.959 1.229
ArtGS liu2025building 1.921 0.677 0.934 0.637 0.506 1.154 1.131 0.502 1.006 0.787 0.926
@@
CD(ws) $\downarrow$ A-SDF mu2021sdf 112.30 34.01 94.11 33.46 56.73 55.82 3.78 27.64 29.60 25.30 47.28
PARIS liu2023paris 2.09 14.51 12.81 2.64 10.42 18.30 0.80 3.25 10.89 5.89 8.16
D-3DGS yang2024deformable 98.90 16.79 91.84 48.78 30.10 30.93 1.262 65.11 22.39 18.79 42.49
ArtGS liu2025building 2.34 1.72 0.68 2.05 0.52 20.59 0.63 1.00 8.78 10.97 4.92
@@
CD(rs) $\downarrow$ A-SDF mu2021sdf 8.992 1.806 1.822 5.166 5.668 7.050 5.991 2.710 5.245 7.794 5.224
PARIS liu2023paris 2.113 0.417 4.704 0.408 1.216 0.857 4.442 1.354 2.793 4.332 2.264
D-3DGS yang2024deformable 9.589 1.544 0.891 0.465 1.147 3.218 0.168 0.449 3.773 3.100 2.434
ArtGS liu2025building 4.924 0.670 0.381 0.471 1.622 2.670 1.455 1.269 3.912 0.562 1.794
@@
F1 $\uparrow$ A-SDF mu2021sdf 0.007 0.034 0.056 0.012 0.021 0.023 0.168 0.001 0.010 0.004 0.034
PARIS liu2023paris 0.221 0.151 0.319 0.091 0.423 0.024 0.421 0.530 0.031 0.032 0.224
D-3DGS yang2024deformable 0.064 0.180 0.125 0.073 0.168 0.195 0.233 0.131 0.022 0.023 0.121
ArtGS liu2025building 0.264 0.255 0.454 0.125 0.460 0.055 0.426 0.292 0.062 0.045 0.244
@@
EMD $\downarrow$ A-SDF mu2021sdf 2.123 0.951 0.955 1.606 1.684 1.878 1.731 1.115 1.620 1.837 1.559
PARIS liu2023paris 2.454 0.457 1.533 0.451 0.780 0.655 1.491 0.770 1.183 1.354 1.113
D-3DGS yang2024deformable 3.224 0.612 0.665 0.481 0.758 1.269 0.291 0.436 1.374 1.206 1.032
ArtGS liu2025building 2.207 0.562 0.412 0.563 0.525 1.171 0.631 0.724 1.297 0.750 0.884